OpenStudy (anonymous):

Use a ratio to compare the values. Simplify the ratio. 12 yd to 20 ft A. 5 : 9 B. 3 : 5 C. 5 : 3 D. 9 : 5

OpenStudy (anonymous):

@HazelLuv99

OpenStudy (kendricklamar2014):

The question says, 12 yards to 20 feet. First let's convert the units. One yard is equal to 3 feet. So if we multiply, \[3 * 12 = 36 ~feet\] Now we will, divide 36 feet to 20 feet and we'll simplify. \[\left(\begin{matrix}36 \\ 20\end{matrix}\right)\] both 36 and 20 are divisible by 4, that means 4 is their common factor. We can simplify by that. \[\left(\begin{matrix}36 \\ 20\end{matrix}\right) = \left(\begin{matrix}4*9 \\ 4*5\end{matrix}\right) = 1 * \left(\begin{matrix}9 \\ 5\end{matrix}\right) = \left(\begin{matrix}9 \\ 5\end{matrix}\right)\] The answer is D. 9:5

OpenStudy (kendricklamar2014):

Does this make sense @bleebleebleed

OpenStudy (anonymous):

kind of

OpenStudy (anonymous):

no @KendrickLamar2014

OpenStudy (kendricklamar2014):

What don't you get?

OpenStudy (anonymous):

how you got 9:5

OpenStudy (kendricklamar2014):

All I did was Simplify

OpenStudy (kendricklamar2014):

\[\left(\begin{matrix}4*9 \\ 4*5\end{matrix}\right) = 4/4 * 9/5 = 9/5\]

OpenStudy (kendricklamar2014):

@bleebleebleed

OpenStudy (anonymous):

ohhh @KendrickLamar2014

OpenStudy (kendricklamar2014):

\[?\]

OpenStudy (anonymous):

? @KendrickLamar2014

OpenStudy (anonymous):

want a medal? @KendrickLamar2014

OpenStudy (kendricklamar2014):

If you want to give me one